OpenStudy (anonymous):

Pyramid Lengths of the sides of its base Height 1 14 cm, 21 cm, and 28 cm 35 cm 2 16 cm, 24 cm and 32 cm 40 cm Which statement is true about the pyramids Chandra drew? Answer They are congruent because the ratio of their corresponding sides is equal. They are similar because the ratio of their corresponding sides is the same. They are not congruent because the ratio of their corresponding sides is equal to one. They are not similar because the ratio of their corresponding sides is not greater than one.

Directrix (directrix):

They are similar because the ratio of their corresponding sides is the same. 14/16=21/24=28/32=35/40